40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 14 00:06:09 2020, gautam, Update, LSC, Locking with POX for CARM 
    Reply  Tue Jul 14 01:55:03 2020, Koji, Update, LSC, Locking with POX for CARM 
    Reply  Tue Jul 14 17:28:29 2020, gautam, Update, LSC, Locking with POX for CARM PRFPMI_2020712sensMat.pdfLSCerrSigs.pdf
Message ID: 15481     Entry time: Tue Jul 14 17:28:29 2020     In reply to: 15476
Author: gautam 
Type: Update 
Category: LSC 
Subject: Locking with POX for CARM 

From Attachment #1, looks like the phasing and gain for CARM on POX11 is nearly the same as CARM of REFL11, which is probably why I was able to execute a partial transition last night. The response in POY11 is ~10 times greater than POX11, as expected - though the two photodiodes have similar RF transimpedance, there is a ZFL-500-HLN at the POY11 output. The actual numerical values are 2.5e10 cts/m for CARM-->REFL11_I, 2.6e10 cts/m for CARM-->POX11_I, and 3.2e11 cts/m for CARM-->POY11_I.

So I think I'll just have to fiddle around with the transition settings a little more tonight. 

One possible concern is that the POX and POY signals are digitized without preamplificatio, maybe this explains the larger uncertainty ellipse for the POX and POY photodiodes relative to the REFL11 photodiode? Maybe the high frequency noise is worse and is injecting junk in the AO path? I think it's valid to directly compare the POX and REFL spectra in Attachment #2, without correcting for any loops, because this signal is digitized from the LSC demodulator board output (not the preamplified one, which is what goes to the CM board, and hence, is suppressed by the CARM loop). Hard to be sure though, because while the heads are supposed to have similar transimpedance, and the POX photodiode has +12dB more whitening gain than REFL11, and I don't know what the relative light levels on these photodiodes are in lock.


I have some data from a couple of days ago when the PRFPMI was locked as usual (CARM_B on REFL for both DC and AO paths), and the sensing lines were on, so I can measure the relative strength of the sensing lines in POX/REFL and get an estimate of what the correct digital gain should be

Attachment 1: PRFPMI_2020712sensMat.pdf  56 kB  Uploaded Tue Jul 14 18:35:10 2020  | Show | Show all
Attachment 2: LSCerrSigs.pdf  120 kB  Uploaded Tue Jul 14 18:55:04 2020  | Show | Show all
ELOG V3.1.3-